Liberty Services is now at the end of the final year of a project. The equipment originally cost $125,500, of which 75% has been depreciated. The firm can sell the used equipment today for $36,000, and its tax rate is 30%. What is the equipment’s after–tax salvage value for use in capital budgeting analysis?
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=125500*(1-0.75)=31375

Hence gain on sale=36,000-31375=$4625

Hence after-tax salvage value=Sale proceeds-(gain on sale*Tax rate)

=36,000-(4625*0.3)

=$34612.5